differentiate the following function using derivative rules:
f(x)= (2+ ln x)^3

Answer by Boreal(15235)

f(x)=(2 + ln (x))^3
f'(x)=3(2+ln x)^2*(1/x), the derivative of what is in the parentheses.
f'(x)=(3/x)(2+ln x)^2